Using a 0.25 M phosphate buffer with a pH of 7.2, you add 0.75 mL of 0.55 M HCl to 57 mL of the buffer. What is the new pH of the solution? (Enter your answer to three significant figures.)


Using a 0.25 M phosphate buffer with a pH of 7.2, you add 0.75 mL of 0.55 M NaOH to 57 mL of the buffer. What is the new pH of the solution? (Enter your answer to three significant figures.)
WebAssign will check your answer for the correct number of significant figures.
Answer #1

The Ka values for the phosphoric acid are The phosphoric acid has the following ionizations H3PO4 = H2PO4 + H+ pka2 = 2.15 H2

(a) Number of moles of HCl added = 0.55 M * 0.75 mL = 0.4125 mmol When HCl adds to the buffer solution, it reacts with HPO-2

(b) Number of moles of NaOH added = M*V = 0.55 M * 0.75 mL = 0.4125 mmol When NaOH adds to the buffer solution, it reacts wit